OpenStudy (anonymous):

If the arc on a particular circle has an arc length of 14 inches, and the circumference of the circle is 84 inches, what is the angle measure of the arc?

OpenStudy (anonymous):

A. 60° B. 90° C. 120° D. 180°

OpenStudy (campbell_st):

well you know the are length and the circumference so comparing the equations, x is the sector angle formed by the arc \[\frac{\frac{\x}{360} \times \pi \times d}{\pi \times d} = \frac{14}{84}\] cancel the common factors shows that \[\frac{\x}{360} = \frac{14}{84}\] solve for x

OpenStudy (anonymous):

is it 60?

OpenStudy (campbell_st):

correct